OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 5.2x + 3.2 = 9.8 A. 1.1 B. 1.4 C. 1.5 D. 1.2

OpenStudy (anonymous):

1. Identify. 5.2x+3.2=9.8 x could be 1.1, 1.4, 1.5, or 1.2 2. Lay it out You will want to lay these all out to see what your equations are. 5.2•1.1+3.2=? 5.2•1.4+3.2=? 5.2•1.5+3.2=? 5.2•1.2+3.2=? 3. Round Since all of the decimal #'s in the equation are below 4, you round down BUT the answer goes up to 10! 5•1.1+3=? 5•1.4+3=? 5•1.5+3=? 5•1.2+3=? 4. Plug it in 5•1.1+3=8.5 5•1.4+3=10 5•1.5+3=10.5 5•1.2+3=9 So your answer is b. I hope this was helpful :3

OpenStudy (anonymous):

@Death_The_Kid Yay! that was the answer I got too and I was checking to make sure I'm correct, thank you! :)

OpenStudy (anonymous):

Your very welcome!